# An App for the Discovery of Properties of Poncelet Triangles

### v8

@@ 1. Introduction
As shown in Figure 3, six concentric, axis-parallel Poncelet triangle families are supported [^1] Another dozen other non-Ponceletian families are supported where two vertices remain stationary while a third one rides on a curve related to a base ellipse., as well as three additional non-concentric ones, Figure 4.
@@
Figure 4. Four circle-inscribed families supported: inellipse (top-left): “inellipse”, caustic is a concentric inellipse (appears in Figure 3); poristic (top-right): Chapple’s porism, caustic is the fixed incircle; macbeath (bottom-left), excentrals to the poristic family, caustic is the MacBeath inconic [40], with foci on $X_{3}$ and $X_{4}$; brocard (bottom-right): also known as the Brocard porism [3], or the “harmonic family”, caustic is the Brocard inellipse [40], with foci on the two Brocard points $\Omega_{1},\Omega_{2}$.
@@
Dual $X_{4}$ Y N n/a
Inellipse $X_{3}$ Y Y $\sum{s_{i}^{2}}$, $\prod\cos\theta_{i}$,$r_{h}$,$R_{h}$,$(\kappa_{i}^{\prime})^{2/3}$
@@
MacBeath $X_{3},X_{4}$ N Y $\sum\cos(2\theta_{i})$, $\prod\cos(\theta_{i})$, $\sum{s_{i}}^{2}$
Table 1. Each of the 9 Poncelet triangle families supported by the app, and some triangle centers which remain fixed. The “CAP” (resp. “CI”) column indicates if the family is interscribed in a concentric-axis parallel pair of ellipses (resp. if the outer conic is a circle). The conserved quantities in the last column are a result of experimental work, and many have been proved. Symbols $L$,$J$,$r$,$R$,$A$, $\omega$ refer to perimeter, Joachmisthal’s constant [39], inradius, circumradius, area, and Brocard angle, respectively. $s_{i}$, $\theta_{i}$, $\kappa_{i}$ refer to sidelength, internal angle, and outer ellipse curvature, respectively, $i=1,...N$, where $N$ is the number of sides of a given Poncelet family. Primed (resp. $h$-subscripteD) symbols refer to quantities measured in the excentral (resp. orthic) triangle.
@@ 4.1. Poncelet families
  • Ellipse-inscribed, Figure 3: (i) Confocal (i.e., elliptic billiard), (ii) Homothetic, (iii) with Incircle, (iv) Dual, (v) Excentral (to confocals).
  • Circle-inscribed, Figure 4: (vi) Poristic ($N=3$ bicentrics), (vii) Brocard Porism, (viii) MacBeath (excentrals to poristic), (ix) Concentric inellipse.
@@ 6.1. Centers and Vertices
    ort: the orthopole of line $X_{m}X_{n}$, see [40, Orthopole], where $m$ and $n$ are selected triangle and cevian centers, see Section 7 and Section 8.
@@ 8.1. Traditional
Available in this groups are the standard constructions for (i) Cevian, (ii) Anticevian, (iii) Circumcevian, (iv) Pedal, (v) Antipedal, and (vi) Trilinear Polar triangles described in [40]. Recall that the latter produces a degenerate (segment-like) triangle, see [40, Trilinear Polar]. Additionally, the setting ccev-simson computes the triangle bounded by the Simson lines corresponding to the vertices of the $X_{m}$-circumcevian triangle.
@@ 8.2. Inversive
  • polar: $T_{m}$ will be bounded by the polars (infinite lines) of the parent’s vertices with respect to a unit circle centered on $X_{m}$, see [40, Polar].
  • inv-excircs: $T_{m}$ will have vertices at inversions of $X_{m}$ with respect to its excircles, see [40, Excircle].
@@ 9.1. Ellipse-Affixed Circles
  • *ctr*: in a pair of concentric ellipses, the common center, else the center of the circumcircle (poristic and brocard).
  • *f1*,*f2*: the left (resp. right) focus of the outer ellipse. Note: in the poristic (resp. excentral) family this becomes the center of the outer circle (resp. caustic = elliptic billiard).
@@ 9.2. Central Circles
Most of these are defined in [40, Central Circles]:
@@
  • gheorghe: Gheorghe’s circle, see [18, X(649)]
@@
  • polar: Polar circle, centered on $X_{4}$. Since radius for acute triangles is negative, we show it with absolute radius) [40, Polar circle].
@@ 11.2. Detection of Metric Invariants
  • $\sum{s^{2}_{i}},\sum{1/s_{i}},\sum{s_{i}^{-2}}$: sum of squared, reciprocal, or reciprocal-squared sidelengths, respectively.
  • $\prod{s_{i}}$: the product of sidelengths.
@@
  • $\sum{\sin},\sum{\cos},\sum{\tan}$: sum of internal angle sines, cosines, and tangents.
  • $\sum{\sin(2t)},\sum{\cos(2t)},\sum{\tan(2t)}$,$\sum{\cot(2t)}$: sum of double-angle trigonometric quantities
  • $\sum{\sin(t/2)},\sum{\cos(t/2)},\sum{\tan(t/2)}$,$\sum{\cot(t/2)}$: sum of half-angle trigonometric quantities.
@@
  • $r_{pol}^{2}$: squared-radius of the triangle’s “polar circle”, $r_{pol}^{2}=R^{2}-(\sum{s_{i}}^{2})/2$, negative for acute triangles [40, Polar Circle].
Also reported whenever a derived triangle is selected, are one of $L^{\prime}/L$, $A^{\prime}/A$, $A^{\prime}.A$, $R_{c}^{\prime}/R_{c}$ if these are invariant.
@@ 16.2. Research
  • 24 poristic tangential and orthic (8 items): locus of the so-named triangles over the poristic family
  • 25 circumcimson & bitangs (11 items): experiments with the circumcimson and various bitangent triangles over the poristic and incircle families

### v9

@@ 1. Introduction
The app supports the following groups of Poncelet families:
  1. (1)
    Figure 3: six ellipse-inscribed families with concentric, axis-parallel caustics.
  2. (2)
    Figure 4: four non-concentric, circle-inscribed families.
  3. (3)
    Figure 5: four ellipse-inscribed, non-concentric, circular caustic families.
An additional dozen-odd non-Ponceletian families are supported where two vertices remain stationary while a third one rides on a curve related to a base ellipse.
@@
Figure 4. Four circle-inscribed families supported: inellipse (top-left): “inellipse”, caustic is a concentric inellipse (appears in Figure 3); bicentric (top-right): Chapple’s porism, caustic is the fixed incircle; macbeath (bottom-left), excentrals to the bicentric family, caustic is the MacBeath inconic [41], with foci on $X_{3}$ and $X_{4}$; brocard (bottom-right): also known as the Brocard porism [3], or the “harmonic family”, caustic is the Brocard inellipse [41], with foci on the two Brocard points $\Omega_{1},\Omega_{2}$.
Figure 5. Four ellipse-inscribed families with non-concentric, circular caustics: top-left: stationary incenter $X_{1}$ (at a focus); top-right: stationary barycenter $X_{2}$ (and Nagel point $X_{8}$ at center); bottom-left: stationary orthocenter $X_{4}$ (at a focus); bottom-right: [to do] stationary Gergonne point $X_{7}$. Caustic centers and radii kindly derived by Ronaldo Garcia [9].
@@
Dual $X_{4}$ Y N $\rho^{2}=4R^{2}-(\sum{s_{i}^{2}})/2$
Inellipse $X_{3}$ Y Y $\sum{s_{i}^{2}},\prod\cos\theta_{i},r_{h},R_{h},(\kappa_{i}^{\prime})^{2/3},\rho^{2}$
@@
MacBeath $X_{3},X_{4}$ N Y $\sum\cos(2\theta_{i}),\prod\cos(\theta_{i}),\sum{s_{i}}^{2},\rho^{2}$
iso-$X_{1}$ $X_{1}$ N N $\sum\sin(\theta_{i}/2)$
iso-$X_{2}$ $X_{k},k=1,2,8,10,...$ N N n/a
iso-$X_{4}$ $X_{k},k=1,4,7952,18283,...$ N N $\rho^{2}$
iso-$X_{7}$ $X_{k},k=1,7,20,77,170,...$ N N $\sum\tan(\theta_{i}/2)$
Table 1. Each of the 9 Poncelet triangle families supported by the app, and some triangle centers which remain fixed. The “CAP” (resp. “CI”) column indicates if the family is interscribed in a concentric-axis parallel pair of ellipses (resp. if the outer conic is a circle). The conserved quantities in the last column are a result of experimental work, and many have been proved. Symbols $L$,$J$,$r$,$R$,$A$, $\omega$ refer to perimeter, Joachmisthal’s constant [40], inradius, circumradius, area, and Brocard angle, respectively. $s_{i}$, $\theta_{i}$, $\kappa_{i}$ refer to sidelength, internal angle, and outer ellipse curvature, respectively, $i=1,...N$, where $N$ is the number of sides of a given Poncelet family. Primed (resp. $h$-subscripteD) symbols refer to quantities measured in the excentral (resp. orthic) triangle. The quantity $\rho^{2}$ is the squared radius of the polar circle, negative for acute triangles [41, Polar Circle].
@@ 4.1. Poncelet families
  • Ellipse-inscribed, concentric and aligned caustic, Figure 3: (i) Confocal (i.e., elliptic billiard), (ii) Homothetic, (iii) with Incircle, (iv) Dual, (v) Excentral (to confocals).
  • Circle-inscribed, Figure 4: (vi) Bicentric, (vii) Brocard Porism, (viii) MacBeath (excentrals to bicentric), (ix) Concentric inellipse.
  • Ellipse-inscribed, non-concentric, circular caustic, Figure 5, where circ-$X_{k}$ implies that $X_{k}$ is stationary:
    1. (1)
      circ-$X_{1}$ (aka. ‘incenter-focal’), tangentials to the bicentric family;
    2. (2)
      circ-$X_{2}$ (aka. ‘iso-baric’);
    3. (3)
      circ-$X_{4}$ (aka. ‘ortho-focal’);
    4. (4)
      circ-$X_{7}$ (aka. ‘iso-Gergonne’), tangentials to the Brocard family.
    5. (5)
      Additionally, circ-Half and circ-Third families are included where the circular caustic is centered at $(a/2,b/2)$ and $(a/3,b/3)$ respectively.
@@ 6.1. Centers and Vertices
    ort: the orthopole of line $X_{m}X_{n}$, see [41, Orthopole], where $m$ and $n$ are selected triangle and cevian centers, see Section 7 and Section 8.
@@ 8.1. Traditional
Available in this groups are the standard constructions for (i) Cevian, (ii) Anticevian, (iii) Circumcevian, (iv) Pedal, (v) Antipedal, and (vi) Trilinear Polar triangles described in [41]. Recall that the latter produces a degenerate (segment-like) triangle, see [41, Trilinear Polar]. Additionally, the setting ccev-simson computes the triangle bounded by the Simson lines corresponding to the vertices of the $X_{m}$-circumcevian triangle.
@@ 8.2. Inversive
  • polar: $T_{m}$ will be bounded by the polars (infinite lines) of the parent’s vertices with respect to a unit circle centered on $X_{m}$, see [41, Polar].
  • inv-excircs: $T_{m}$ will have vertices at inversions of $X_{m}$ with respect to its excircles, see [41, Excircle].
@@ 9.1. Ellipse-Affixed Circles
  • *ctr*: in a pair of concentric ellipses, the common center, else the center of the circumcircle (bicentric and Brocard).
  • *f1*,*f2*: the left (resp. right) focus of the outer ellipse. Note: in the bicentric (resp. excentral) family this becomes the center of the outer circle (resp. caustic = elliptic billiard).
@@ 9.2. Central Circles
Most of these are defined in [41, Central Circles]:
@@
  • gheorghe: Gheorghe’s circle, see [19, X(649)]
@@
  • polar: Polar circle, centered on $X_{4}$. Since radius for acute triangles is negative, we show it with absolute radius) [41, Polar circle].
@@ 11.2. Detection of Metric Invariants
  • $\sum{s_{i}^{-2}},\sum{s^{2}_{i}},\sum{s_{i}^{-2}},\sum{\sqrt{s_{i}}},\sum{1/\sqrt{s_{i}}}$: sum of inverse, squared, inverse-squared, square-root, and inverse-square-root sidelengths, respectively.
  • $\sum{s_{ij}},\sum{s_{ij}^{-2}},\sum{s^{2}_{ij}},\sum{s_{ij}^{-2}}$: sum of plain, inverse, squared, and inverse-squared ‘cyclic’ sidelength products, $\{s_{ij}\}=\{s_{1}s_{2},s_{2}s_{3},s_{3}s_{1}\}$, respectively.
  • $\prod{s_{i}}$: product of sidelengths.
@@
  • $\sum{\sin},\sum{\cos},\sum{\tan},\sum\sec,\sum\csc$: sum of internal angle sines, cosines, tangents, secants, and co-secants.
  • $\sum{\sin(2t)},\sum{\cos(2t)},\sum{\tan(2t)},\sum{\sec(2t)},\sum{\csc(2t)},\sum{\cot(2t)}$: sum of double-angle trigonometric quantities
  • $\sum{\sin(t/2)},\sum{\cos(t/2)},\sum{\tan(t/2)},\sum{\sec(t/2)},\sum{\csc(t/2)},\sum{\cot(t/2)}$: sum of half-angle trigonometric quantities.
@@
  • $r_{pol}^{2}$: squared-radius of the triangle’s “polar circle”, $r_{pol}^{2}=4R^{2}-(\sum{s_{i}}^{2})/2$, negative for acute triangles [41, Polar Circle].
Also reported:
  • When a derived triangle is selected: invariant $L^{\prime}/L$, $A^{\prime}/A$, $A^{\prime}.A$, $R/R^{\prime}$ ratios.
  • When a central circle is selected: invariant radius $R_{c}$.
  • When both a derived triangle and central circle are selected: $R_{c}^{\prime}/R_{c}$ is constant.
@@ 16.2. Research
  • 24 bicentric tangential and orthic (8 items): locus of the so-named triangles over the bicentric family
  • 25 circumcimson & bitangs (11 items): experiments with the circumcimson and various bitangent triangles over the bicentric and incircle families